OpenStudy (anonymous):

Which of the following equations represents the graph of a line that is perpendicular to the graph of y = kx + b (where k and b are constants) and goes through the point (3, -4)? A) y + 4 = k(x - 3) B) y - 4 = k(x + 3) C) y + 4 = (-1/k)(x - 3) D) y - 4 = (-1/k)(x + 3)

OpenStudy (anonymous):

1. Find the slope of the perpendicular, using slope1*slope2 = -1 2. Find the intercept using the given points

OpenStudy (anonymous):

Use the imformation given: Point (3,-4) =\[(x_{1},y_{1})\] In order to be perpendicular it slope must be the reciprocal negative of the other line that is m = -1/k Use the slope point formula : \[y-y_{1}=m(x-x_{1})\] plug in the information given in the formula \[y +4 = \frac{ -1 }{ k}\left( x-3)\right)\]